# references.bib

### v1

@@ 6.3.2 Training AI in the EU: Ongoing Controversies
Conditions for processing user data to train AI have been a contentious topic in the European Union. Throughout 2024, tech companies like Google, Meta, X and LinkedIn have delayed or suspended plans to launch generative AI models in the EU due to regulatory inquiries into their training practices [politico]. So far, these companies have been aiming to harness user-generated content to train their AI assistants or large language models and have not been openly targeting private messaging. Nevertheless, AI training practices have already been raising substantial privacy concerns in the EU.
@@
Ireland’s Data Protection Commissioner (DPC) has actively monitored the release of several AI-powered virtual assistants fueled by EU user data. The Irish authority is the European Union’s lead regulator for many US tech companies due to their EU headquarters being based in Dublin. In June 2024, the DPC raised concerns about Meta’s privacy policy update that allowed Meta AI to be trained on public posts of EU users on Facebook and Instagram [dpcmeta]. The DPC acted in response to a complaint filed by the advocacy group NOYB - European Center for Digital Rights, founded by privacy activist Max Schrems. NOYB filed complaints in eleven countries urging regulators to act against Meta’s “abuse of personal data for AI”, drawing attention to the feature’s lack of opt-in consent and hidden opt-out form [noybmetaai].
@@
Similarly, in August 2024, the DPC initiated proceedings against X (formerly Twitter) in the Irish High Court for concerns related to the training of X’s AI assistant Grok [dpcgrok]. A few months prior, users noticed a default setting on the app, collecting user consent via a pre-ticked box, allowing X to use posts, interactions, inputs and results for AI-training [guardiangrok]. The proceeding was dismissed after X agreed to suspend the use of publicly posted data on X to train Grok. In September, the DPC opened an inquiry into Google’s use of European Union users’ personal data to help develop its foundational AI Model, Pathways Language Model 2 (PaLM 2) [reuterspalm]. The DPC also publicly confirmed that it has been in contact with Apple regarding its plans for Apple Intelligence AI [ieapple]. At the broader industry level, an open letter coordinated by Meta and signed by executives from more than two dozen companies has criticized the EU’s regulatory approach for generating uncertainty and for blocking innovation in the region [wsjmeta] [metaopenletter].
@@
Regulator inquiries and guidelines from EU member states have typically operated on a case-by-case basis, taking into consideration each AI system’s business model and operational specifications. The European Data Protection Board (EDPB), which oversees the consistent application of data protection rules across the EU, began harmonizing the GDPR’s interpretation regarding AI training this year. Following the launch of ChatGPT and the Italian regulator’s enforcement action against OpenAI, the EDPB launched a dedicated task force to investigate privacy issues in the ChatGPT service [edpbchatgpt]. The main findings for AI training were based on OpenAI’s practice of “web scraping” (the automatic extraction of data from third-party platforms). This practice is different from obtaining data posted by users on proprietary platforms, as is currently the case in MetaAI and Grok’s processing operations. Nevertheless, some of the EDPB’s recommendations may be more broadly applicable, such as the requirement that users be clearly and demonstrably informed that any prompts, including text queries and file uploads, may be used for training purposes.
@@
Aiming to trigger more comprehensive discussions at the supranational level, the DPC requested the EDPB for a supervisory opinion on relevant issues when processing personal data to develop and train AI models [dpcgrok]. The EDPB’s much-awaited Opinion 48/2024 was published in December 2024, focusing specifically on (1) when and how an AI model can be considered as “anonymous”; (2) how organizations can demonstrate the appropriateness of legitimate interest as a legal basis in the development and (3) deployment phases; and (4) consequences of the unlawful processing of personal data in the development phase of an AI model on subsequent processing operations [edpbdecai]. With regard to balancing organizations’ “legitimate interests” and user rights, the EDPB highlighted the importance of transparency and meeting users’ reasonable expectations considering the complexity of the technologies and the variety of potential uses of personal data. The EDPB did not address user consent; the EDPB noted that the opinion does not cover all EU data protection issues in the context of AI development and deployment.
@@
Importantly, while the EU is openly debating the lawfulness of AI training under different conditions, none of the regulator inquiries or opinions to date address the issue of using data obtained from messaging or end-to-end encrypted services. While data posted on social media platforms under public settings may generally be considered public, messaging data is private by nature. [^48] There are also nuances regarding the public nature of publicly posted data: “(…) the mere fact that personal data is publicly accessible does not imply that ’the data subject has manifestly made such data public’. In order to rely on the exception laid down in Article 9(2)(e) GDPR, it is important to ascertain whether the data subject had intended, explicitly and by a clear affirmative action, to make the personal data in question accessible to the general public.” See [edpbchatgpt]. Legal regimes typically recognize that private communications must be treated to a higher standard of privacy protection because of their more intimate and vulnerable qualities, as embodied by the EU’s e-Privacy directive. [^49] In the US, the “Federal Wiretap Act” (18 U.S.C. §§ 2510-2522), also known as “Title III”, prohibits the interception and disclosure of wire, oral and electronic communications. Thus, the use of private messages shifts the paradigm and necessarily calls for more comprehensive debates around user consent. Moreover, training on what is advertised or commonly thought to be E2EE data may frustrate users’ legitimate expectations as to the privacy and security of their data.
@@
Although the regulation of AI training in the EU is still in flux and conclusions are uncertain, the recent inquiries in the EU and the UK have indicated that simply updating the company’s privacy policy is insufficient to comply with GDPR standards. Besides providing users and regulators with proactive transparency and disclosure, the substance of how companies are collecting personal data, the types of data harnessed and how they are being used also matter. If private messages are used as training data, questions related to the indispensability of user consent, adequate interfaces for collecting consent and options to exercise control over data become substantially more pressing.
@@ 6.5 Privacy, Antitrust and Consumer Data
At the time of Meta’s acquisition of WhatsApp in 2014, the European Commission approved the acquisition [ecmerger1] with the understanding that Facebook would be unable to do automated matching between Facebook users’ accounts and WhatsApp users’ accounts [ecmerger2]. However, in August 2016 WhatsApp announced updates to its terms of service and privacy policy including the possibility of linking user data across WhatsApp and Facebook. The aftermath resulted in a €110 million fine on Meta (formerly Facebook) for providing misleading information under EU merger regulation [ecmerger2]. Had a separation between WhatsApp and Facebook data been enforced beyond levying a fine, this antitrust measure could have had far-reaching implications for the integration of AI assistants across WhatsApp and Facebook.
In November 2024, India’s antitrust regulator targeted Meta for anti-competitive behavior related to data sharing [indiaexpress]. The Competition Commission of India (CCI) imposed a penalty of roughly 25.4 million dollars on Meta after WhatsApp users were forced to to share their data with other Meta-owned companies by means of a 2021 privacy policy update with no opt-out mechanism [indiareuters]. The regulator also ordered WhatsApp to refrain from sharing user data for advertising purposes with other Meta-owned applications for five years, highlighting that the data sharing resulted in denial of market access for competitors.
These examples illustrate how data privacy decisions can lead to broader antitrust concerns beyond the scope of privacy and vice versa. While companies often have monetary incentives to access larger pools of data and integrate features throughout a broad range of products, this may come at the expense of higher scrutiny by regulators. This is particularly timely observation as the past four years have seen a wave of antitrust lawsuits against Big Tech initiated by the US and the EU [harvardantitrust].
@@ 7 Recommendations
Next, we present our recommendations for integrating AI assistants into E2EE applications. Our recommendations address both compatibility with E2EE from a security perspective (Recommendations 1 and 2, in Section 7.1), and disclosure and consent from a legal and normative perspective (Recommendations 3 and 4, in Section 7.2). As a recap,the summary of our recommendations is as follows: [^50] This bullet-point summary of our recommendations is identical to that in Section 1.
@@
      Prioritize endpoint-local processing whenever possible.
@@
        No third party can see or use any E2EE content [^51] Or any derivative of E2EE content (i.e., any non-trivial function of content). See Section 3.1 for full definitions. without breaking encryption, [^52] This may sound paradoxical; see § 4.2 for detailed discussion of technical configurations that may achieve this condition. and
@@ 7.1 Compatibility with E2EE Security
Next, we describe our recommendations to preserve compatibility with E2EE from a security perspective, informed by our analysis of from Section 4 and guided by our E2EE feature evaluation framework described in the Appendix A.
@@
    Prioritize endpoint-local processing whenever possible.
@@
      No third party can see or use any E2EE content [^54] Or any derivative of E2EE content (i.e., any non-trivial function of content). See Section 3.1 for full definitions. without breaking encryption, [^55] This may sound paradoxical; see § 4.2 for detailed discussion of technical configurations that may achieve this condition. and
@@
Application developers should try to prioritize local inference whenever possible, as this option is fully compatible with E2EE.
If performing non-endpoint-local inference, in order to maintain E2EE confidentiality, it is critical that no third party (including the model provider) handles any E2EE content unencrypted. The one cryptographic tool that might permit this, FHE, is not currently practical for this use (see Section 4.2).
Another approach to protecting the confidentiality of user content when processing it for AI features would be to use TEEs, which are much more practical than FHE. As discussed in Section 2.4.2, TEE security is different from encryption security: a system that relies on TEE security for message confidentiality does not achieve E2EE security and cannot accurately be described as offering E2EE. If exploring TEE-based approaches, we recommend that service providers do not describe these as E2EE (see also Recommendation 3), and exercise great care around user expectations, consent, and potential unanticipated harms if transitioning a system from one type of security guarantee to another (see also Recommendation 4).
@@ 7.2 Towards Meaningful Disclosure and Consent
Consent, then, encompasses assent to the formal terms of service and privacy policies that bind the user and service in a contractual agreement. Opt-in consent mechanisms are those which allow users to take an affirmative step to agree to legal terms, such as pop-up banners for clicking a button or checking a box. These mechanisms may display brief disclosures informing users of the most important terms and may link to the pages where the complete terms and policies are displayed. Conversely, opt-out mechanisms are interfaces that provide consumers with a streamlined process to opt out of certain data processing activities after consent has been given, or as an alternative to the initial opting-in. Opt-out can therefore be considered withdrawal of consent. While users may not read all the fine print (see Section 5.1), the content contained in terms of service meaningfully affects users’ rights and are subject to legally defined floors.
Using these definitions, the line between disclosure and consent can be blurry in some situations. We include terms of service and privacy policies are certainly included under the scope of consent and not disclosure. Elements of a service’s chat interface are included in disclosure, but not necessarily consent. We treat advertising, which is generally not considered binding on a service, as a disclosure. Cookie banners, ubiquitous since the passage of GDPR, overlap both disclosure and consent. Although easily dismissed by many consumers, they are prominently displayed on websites and contain non-legal language designed to be legible, a common characteristic of disclosures. Consumers then select their cookie preferences, making a two-way communication characteristic of consent.
@@
“Unqualified representation” is a term often used in consumer protection law. It generally refers to a statement made without caveats, and, in this paper, excludes caveats made in fine print, such as footnotes or terms of service. So, a splash page for an E2EE messaging service which boldly states “E2EE Used for All Messages*” would be making an unqualified representation, even if the asterisk points to a footnote that says “Unless opted out by user.” or “Except when AI features are used.” Recommendation 3 says that messaging services should not make unqualified claims that they are using E2EE when their implementation does not achieve E2EE confidentiality.
@@
For example, in WhatsApp, many points of the advertising and app interface indicate that “text and voice messages” are E2EE. Many users may see messages invoking Meta AI, which are not encrypted, as a “text or voice message,” and thus expect that their requests to Meta AI are E2EE. Meta reinforces this false, but natural, misunderstanding. When entering a Meta AI chat for the first time, WhatsApp users are shown a screen which states “Personal messages stay private. Meta may use your AI messages to improve AI quality. Personal messages are never sent to Meta, can’t be read, and remain end-to-end encrypted.” The scope of “personal messages” is critical, and ambiguous, in this disclosure. A sufficient disclosure would explicitly state that the messages in the following conversation are not encrypted and are sent to Meta, and as such should not include sensitive information.
@@
Despite normative critiques highlighting the insufficiency of consent, consent is still the main legal tool currently available for giving users greater control over their privacy rights, and implementing consent interfaces well can make an important difference.
@@
We recommend that opt-in express consent be the standard mechanism for allowing messaging services to train AI on user data. Consent for using messaging services must not be conflated with consent for additional features such as AI assistants. Consent for separate features should be collected in separate instances, each with a specifically tailored opt-in banner or other form of disclosure. Although the banner should refer to a more detailed privacy policy and/or terms of service, there should be enough information on the banner to meet previously mentioned disclosure requirements.
@@
Messaging services should recognize the inherently private nature of communications, even (or especially) in environments that mix E2EE and non-E2EE data. Opt-in consent banners should serve the dual purpose of (i) clearly informing users about the exact scope of data that will be used for training purposes, including whether the data is protected by E2EE or not (or by some other security mechanism such as a TEE); (ii) providing users with a real choice, making sure to disclose how their choices may affect E2EE protection. It should be made clear to users which data is kept under E2EE (or TEE) protection, including any differences between individual and group settings. Additionally, terms of service and privacy or security policies should accurately describe all technical security implementations and associated risks.
@@
Digital services should also provide clear opt-out mechanisms upon signing up, including a clear disclosure of what features still work if a user opts out. Each option should carry its own opt-out mechanism. A clear opt-out mechanism is one that is not concealed by dark patterns or unnecessary obstructions. [^56] For instance, requiring users to navigate an eight-step process to access the opt-out form is not a clear mechanism [fpfbrmeta]. As a rule of thumb, opting out or withdrawing consent for services to which users have previously accepted must be made as easily accessible as the initial conditions of consent for using the feature. When possible, opting out from allowing AI training on one’s data should not be a precondition to the use of an entire app or to the use of AI features in the app. Relatedly, services should consider adding separating the opt-in/opt-out for use of an AI feature and for training based on such use. Especially regarding AI training, it must be clearly disclosed that the opt-out is forward-looking: data which the company has already used for training will remain in their models, regardless of any opt-out exercised.575757 Finally, consent should be subject to periodical renewals and opt-in banners should be re-validated after reasonable periods of time.

### v2

@@ 6.3.2 Training AI in the EU and the UK: Ongoing Controversies
Conditions for processing user data to train AI have been a contentious topic in the European Union and the United Kingdom—even for user data in technically public posts, let alone in more private contexts. Throughout 2024, tech companies like Google, Meta, X and LinkedIn have delayed or suspended plans to launch generative AI models in the EU due to regulatory inquiries into their training practices [politico]. So far, these companies have been aiming to harness user-generated content to train their AI assistants or large language models and have not been openly targeting private messaging. Nevertheless, AI training practices have already been raising substantial privacy concerns in the EU.
@@
Ireland’s Data Protection Commissioner (DPC) has actively monitored the release of several AI-powered virtual assistants fueled by EU user data. The Irish authority is the European Union’s lead regulator for many US tech companies due to their EU headquarters being based in Dublin. In June 2024, the DPC raised concerns about Meta’s privacy policy update that allowed Meta AI to be trained on public posts of EU users on Facebook and Instagram [dpcmeta]. The DPC acted in response to a complaint filed by the advocacy group NOYB - European Center for Digital Rights. NOYB filed complaints in eleven countries urging regulators to act against Meta’s “abuse of personal data for AI”, drawing attention to the feature’s lack of opt-in consent and hidden opt-out form [noybmetaai].
@@
Similarly, in August 2024, the DPC initiated proceedings against X (formerly Twitter) in the Irish High Court for concerns related to the training of X’s AI assistant Grok [dpcgrok]. A few months prior, users noticed a default setting on the app, collecting user consent via a pre-ticked box, allowing X to use posts, interactions, inputs and results for AI training [guardiangrok]. The proceeding was dismissed after X agreed to suspend the use of publicly posted data on X to train Grok. In September, the DPC opened an inquiry into Google’s use of European Union users’ personal data to help develop its foundational AI Model, Pathways Language Model 2 (PaLM 2) [reuterspalm]. The DPC also publicly confirmed that it has been in contact with Apple regarding its plans for Apple Intelligence AI [ieapple]. At the broader industry level, an open letter coordinated by Meta and signed by executives from more than two dozen companies has criticized the EU’s regulatory approach for generating uncertainty and for blocking innovation in the region [wsjmeta] [metaopenletter].
@@
Regulator inquiries and guidelines from EU member states have typically operated on a case-by-case basis, taking into consideration each AI system’s business model and operational specifications. The European Data Protection Board (EDPB), which oversees the consistent application of data protection rules across the EU, began harmonizing the GDPR’s interpretation regarding AI training after the launch of ChatGPT. Noting that the Italian regulator took enforcement action against OpenAI, Inc. due to several GDPR violations, the EDPB launched a dedicated task force to investigate privacy issues in the ChatGPT service [edpbchatgpt]. The main findings concerning AI training were based on OpenAI’s practice of web scraping (the automatic extraction of data from third-party platforms). This practice is different from obtaining data posted by users on AI companies’ own platforms, as Meta AI and Grok’s processing operations currently do. Nevertheless, some of the EDPB’s recommendations may be more broadly applicable beyond the specific case of OpenAI and ChatGPT, such as the requirement that users be clearly and demonstrably informed that any prompts, including text queries and file uploads, may be used for training purposes.
@@
Aiming to trigger more comprehensive discussions at the supranational level, the DPC requested the EDPB for a supervisory opinion on relevant issues when processing personal data to develop and train AI models [dpcgrok]. The EDPB’s much-awaited Opinion 28/2024 was published in December 2024, focusing specifically on (1) when and how an AI model can be considered as “anonymous”; (2) how organizations can demonstrate the appropriateness of legitimate interest as a legal basis in the development and (3) deployment phases; and (4) consequences of the unlawful processing of personal data in the development phase of an AI model on subsequent processing operations [edpbdecai]. With regard to balancing organizations’ “legitimate interests” and user rights, the EDPB highlighted the importance of transparency and meeting users’ reasonable expectations considering the complexity of the technologies and the variety of potential uses of personal data. The EDPB did not address user consent; the EDPB noted that the opinion does not cover all EU data protection issues in the context of AI development and deployment.
@@
Importantly, while the EU is openly debating the lawfulness of AI training under different conditions, none of the regulator inquiries or opinions to date address the issue of using data obtained from messaging or end-to-end encrypted services. While data posted on social media platforms under public settings may generally be considered public, messaging data is private by nature. [^58] There are also nuances regarding the public nature of publicly posted data: “(…) the mere fact that personal data is publicly accessible does not imply that ’the data subject has manifestly made such data public’. In order to rely on the exception laid down in Article 9(2)(e) GDPR, it is important to ascertain whether the data subject had intended, explicitly and by a clear affirmative action, to make the personal data in question accessible to the general public.” See [edpbchatgpt]. Legal regimes typically recognize that private communications must be treated to a higher standard of privacy protection because of their more intimate and vulnerable qualities, as embodied by the EU’s ePrivacy directive. [^59] In the US, the Electronic Communications Privacy Act of 1986 protects wire, oral, and electronic communications from unauthorized interception, disclosure, or use, including email, phone calls, and data stored electronically. Thus, the use of private messages shifts the paradigm and necessarily calls for more comprehensive debates around user consent. Moreover, training on what is advertised or commonly thought to be E2EE data may frustrate users’ legitimate expectations as to the privacy and security of their data.
@@
Although the regulation of AI training in the EU is still in flux and conclusions are uncertain, the recent inquiries in the EU and the UK have indicated that simply updating the company’s privacy policy to allow AI training is insufficient to comply with GDPR standards. Besides providing users and regulators with proactive transparency and disclosure, the substance of how companies are collecting personal data, the types of data harnessed and how they are being used also matter. If private messages are used as training data, questions related to the indispensability of user consent, adequate interfaces for collecting consent and options to exercise control over data become substantially more pressing.
@@ 6.5 Privacy, Antitrust and Consumer Data
At the time of Meta’s acquisition of WhatsApp in 2014, the European Commission approved the acquisition [ecmerger1] with the understanding that Facebook would be unable to do automated matching between Facebook users’ accounts and WhatsApp users’ accounts [ecmerger2]. However, in August 2016 WhatsApp announced updates to its terms of service and privacy policy including the possibility of linking user data across WhatsApp and Facebook. The aftermath resulted in a €110 million fine on Meta (formerly Facebook) for providing misleading information under EU merger regulation [ecmerger2]. Had a separation between WhatsApp and Facebook data been enforced beyond levying a fine, this antitrust measure could have had far-reaching implications for the integration of certain operations across WhatsApp and Facebook.
The EU’s Digital Markets Act (DMA), applicable since May 2023, is a landmark antitrust regulation aimed at curbing Big Tech’s market dominance and ensuring fair competition between digital services. The DMA recognizes privacy and data protection as key principles and specifically regulates the prevention of cross-service data merging in Article 5(2). [^60] Regulation (EU) 2022/1925 (”Digital Markets Act - DMA”), Article 5(2). This provision states that gatekeepers cannot combine personal data from different services unless they obtain explicit opt-in consent from users.
Outside the EU, in November 2024, India’s antitrust regulator targeted Meta for anti-competitive behavior related to data sharing [indiaexpress]. The Competition Commission of India (CCI) imposed a penalty of roughly 25.4 million dollars on Meta after WhatsApp users were forced to to share their data with other Meta-owned companies by means of a 2021 privacy policy update with no opt-out mechanism [indiareuters]. The regulator also ordered WhatsApp not to share user data for advertising purposes with other Meta-owned applications for five years, highlighting that the data sharing resulted in denial of market access for competitors.
These examples illustrate how data privacy decisions can lead to broader antitrust concerns beyond the scope of privacy and vice versa. While companies often have monetary incentives to access larger pools of data and integrate features throughout a broad range of products, this may come at the expense of higher scrutiny by regulators. This is a particularly timely observation as the past four years have seen a wave of antitrust lawsuits against Big Tech initiated by the US and the EU [harvardantitrust]. That said, the 2025 government administration changes in the US may change priorities at the FTC, which is a prominent enforcer of both antitrust and privacy laws.
@@ 7 Recommendations
This section presents our recommendations for integrating AI assistants into E2EE applications.
Our recommendations address both compatibility with E2EE from a security perspective (Recommendations 1 and 2, discussed in Section 7.1), and disclosure and consent from a legal and normative perspective (Recommendations 3 and 4, discussed in Section 7.2). We conclude the section with an example application of our recommendations to the practical example of Apple Intelligence (Section 7.3).
As a recap, the bullet-point summary of our recommendations is as follows. [^61] This bullet-point summary of our recommendations is identical to that in Section 1.
@@
      Prioritize endpoint-local processing where possible.
@@
        No third party can see or use any E2EE content [^62] Or any derivative of E2EE content (i.e., any non-trivial function of content). See Section 3.1 for full definitions. without breaking encryption, [^63] This may sound paradoxical; see Section 4.2 for detailed discussion of technical configurations that may achieve this condition. and
@@ 7.1 Compatibility with E2EE Security
    Prioritize endpoint-local processing where possible.
@@
      No third party can see or use any E2EE content [^65] Or any derivative of E2EE content (i.e., any non-trivial function of content). See Section 3.1 for full definitions. without breaking encryption, [^66] This may sound paradoxical; see Section 4.2 for detailed discussion of technical configurations that may achieve this condition. and
@@
Application developers should prioritize local inference whenever possible, as this option is fully E2EE-compatible. To maintain E2EE confidentiality in non-endpoint-local inference, it is critical that no third party (including the model provider) handles any E2EE data unencrypted. [^67] The one cryptographic tool that might permit encrypted processing, FHE, is not currently practical for this use case. See Section 4.2 for more detail.
If exploring TEE-based approaches, service providers should not describe these as E2EE (see also Recommendation 3), and exercise great care around user expectations, consent, and unanticipated harms around transitioning a system from one type of security guarantee to another (see also Recommendation 4). [^68] As discussed in Section 2.4.2, TEE security is different from encryption security: a system that relies on TEE security for message confidentiality does not achieve E2EE security and cannot accurately be described as offering E2EE.
@@ 7.2 Towards Meaningful Disclosure and Consent
Consent, then, encompasses assent to the formal terms of service and privacy policies that bind the user and service in a contractual agreement. Opt-in consent mechanisms are those which allow users to take an affirmative step to agree to legal terms, such as pop-up banners for clicking a button or checking a box. These mechanisms may display brief disclosures informing users of the most important terms and may link to the pages where the complete terms and policies are displayed. Conversely, opt-out mechanisms are interfaces that provide consumers with a streamlined process to opt out of certain data processing activities after consent has been given, or as an alternative to the initial opting-in. Opt-out can therefore be considered withdrawal of consent. While users may not read all the fine print (see Section 5.1), the content contained in terms of service meaningfully affects users’ rights. [^69] Using these definitions, the line between disclosure and consent can be blurry in some situations. We include terms of service and privacy policies under the scope of consent and not disclosure. Elements of a service’s chat interface, like pop-up notifications, are included in disclosure, but not necessarily consent. We treat advertising, which is generally not considered binding on a service, as a disclosure. Cookie banners, ubiquitous since the passage of GDPR, overlap both disclosure and consent: although easily dismissed by many consumers, they are prominently displayed on websites and contain non-legal language designed to be legible, a common characteristic of disclosures. Consumers then select their cookie preferences, making a two-way communication characteristic of consent.
@@
“Unqualified representation” is a term often used in consumer protection law. It generally refers to a statement made without caveats, and, in this paper, excludes caveats made in fine print, such as footnotes or terms of service. So, a splash page for an E2EE messaging service which boldly states “E2EE Used for All Messages*” would be making an unqualified representation, even if the asterisk points to a footnote that says “Unless opted out by user.” or “Except when AI features are used.”
Put more simply, Recommendation 3 says that messaging services should not make unqualified claims that they are using E2EE when their implementation does not achieve E2EE confidentiality.
@@
Figure 6: WhatsApp’s disclosure upon a user’s first use of Meta AI as of 15 November 2024
For example, in WhatsApp, many points of the advertising and app interface indicate that “text and voice messages” are E2EE. Many users may see messages invoking Meta AI, which are not encrypted, as a “text or voice message,” and thus expect that their requests to Meta AI are E2EE. Meta reinforces this false, but natural, misunderstanding. As shown in Figure 6, when entering a Meta AI chat for the first time, WhatsApp users are shown a screen which states “Personal messages stay private. Meta may use your AI messages to improve AI quality. Personal messages are never sent to Meta, can’t be read, and remain end-to-end encrypted.” The scope of “personal messages” is critical, and ambiguous, in this disclosure. A sufficient disclosure would explicitly state that the messages in the following conversation are not encrypted and are sent to Meta, and as such should not include sensitive information.
@@
Despite normative critiques highlighting the insufficiency of consent (see Section 5.2.2), consent is still the main legal tool currently available for giving users greater control over their privacy rights. Implementing consent interfaces well can make an important difference.
@@
We recommend that opt-in consent be the standard mechanism for allowing messaging services to train AI on user data. Consent for using messaging services must not be conflated with consent for additional features such as AI assistants. Consent for separate features should be collected in separate instances, each with a specifically tailored opt-in banner or other form of disclosure. Although the banner should refer to a more detailed privacy policy and/or terms of service, there should be enough information on the banner to meet previously mentioned disclosure requirements.
@@
Messaging services should recognize the inherently private nature of communications, even (or especially) in environments that mix E2EE and non-E2EE data. Opt-in consent banners should serve the dual purpose of (i) clearly informing users about the exact scope of data that will be used for training purposes, including whether the data is protected by E2EE or not (or by some other security mechanism such as a TEE); (ii) providing users with a real choice, making sure to disclose how their choices may affect E2EE protection. It should be made clear to users which data is kept under E2EE (or TEE) protection, including any differences between individual and group settings. Additionally, terms of service and privacy or security policies should accurately and adequately describe security-relevant implementation details and associated risks of enabling the AI assistant feature.
Beyond legal considerations, opt-in consent for AI features better preserves the application’s systemic E2EE security properties (Section 2.1.3): user inaction results in the unaltered functioning of E2EE protections, while only users who explicitly want the E2EE-violating feature will choose to have it enabled.
@@
Finally, digital services should also provide clear opt-out mechanisms upon signing up and when new features are added, including a clear disclosure of what features still work if a user opts out. Each option should carry its own opt-out mechanism. A clear opt-out mechanism is one that is not concealed by dark patterns or unnecessary obstructions. [^70] For instance, requiring users to navigate an eight-step process to access the opt-out form is not a clear mechanism [fpfbrmeta]. As a rule of thumb, opting out or withdrawing consent for services to which users have previously accepted must be made as easily accessible as the initial consent for using the feature. When possible, opting out from allowing AI training on one’s data should not be a precondition to the use of an entire app or to the use of AI features in the app. Relatedly, services should consider separating the opt-in/opt-out for use of an AI feature and for training based on such use. Especially regarding AI training, it must be clearly disclosed that the opt-out is forward-looking: data which the company has already used for training will remain in their models, regardless of any opt-out exercised. [^71] Implementing retroactive withdrawal is currently practically infeasible [goel2024corrective]. Finally, consent should be subject to periodic renewals and opt-in banners should be re-validated after reasonable periods of time.